# Learning tensor networks with parameter dependence for Fourier-based option pricing

### v1

@@ V.2.1 The case of varying $\vec{\sigma}$
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec]
$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $5.63\times 10^{4}$ $5.0\times 10^{5}$ $6.58\times 10^{-4}$ $2.90\times 10^{-2}$
$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $1.16\times 10^{5}$ $6.0\times 10^{5}$ $7.92\times 10^{-4}$ $3.28\times 10^{-2}$
$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $1.78\times 10^{5}$ $7.0\times 10^{5}$ $1.15\times 10^{-3}$ $3.66\times 10^{-2}$
$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0258 0.0340 $3.06\times 10^{5}$ $8.0\times 10^{5}$ $1.67\times 10^{-3}$ $3.93\times 10^{-2}$
$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $4.27\times 10^{5}$ $9.0\times 10^{5}$ $2.98\times 10^{-3}$ $4.54\times 10^{-2}$
$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $6.36\times 10^{5}$ $1.0\times 10^{6}$ $4.47\times 10^{-3}$ $4.93\times 10^{-2}$
$11$ 0.48 $6.0\times 10^{-6}$ $6.0\times 10^{-6}$ 0.0389 0.0202 $6.38\times 10^{5}$ $1.1\times 10^{6}$ $4.22\times 10^{-3}$ $4.88\times 10^{-2}$
$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0377 0.0401 $4.83\times 10^{6}$ $1.1\times 10^{6}$ $1.73\times 10^{-2}$ $5.28\times 10^{-2}$
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{i}\in[0.15,0.25]$ and $S_{0}^{i}\in[90,120]$, and place 100 equally spaced grid points within these ranges. The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$. For both (a) $\vec{\sigma},d=10$ and (b) $\vec{S}_{0},d=11$, there are two scenarios for $\eta$: constant and auto. The term auto in (a) and (b) indicate that $\eta_{i}$ was automatically determined for each asset based on Eqs. (17) and (18), respectively. In the constant scenario, the maximum relative error does not stay within the error range of MC with PATH$=10^{5}$, whereas the auto scenario achieves success. $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs. In TT-based option pricing, the sharp increase in computational complexity around $d=10$ arises. This stems from the necessity to set $\eta$ appropriately for each asset, which in turn makes the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ for $\vec{p}$.
@@ V.2.2 The case of varying $\vec{S_{0}}$
Figure 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. Up to $d=10$, TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. At $d$=5, after compression using SVD, the bond dimensions were reduced to around 10 (refer to Fig. 3(b)). However, at $d=11$, due to the same reasons as in the case of $\sigma$, the computational complexity of TT-based calculations becomes comparable to that of the Monte Carlo method. Nonetheless, due to the different units of computational complexity, we saw superiority over the Monte Carlo method in terms of computation time. Compared to varying $\sigma$, its accuracy is worse at $d=10$.
@@ V.2.3 Randomness in learning the TTs
Here, we mention the randomness in learning the TTs and the error induced by it. Since TCI starts with the randomly selected grid points, the accuracy of the TT-based method depends on these initial points. To assess such a fluctuation of the accuracy, for $d=10$, we evaluated the mean and standard deviation of the relative accuracy of the TT-based method in 20 runs, with initial points randomly selected in each. As a result, the mean was 0.0324 and the standard deviation was 0.0039, which indicates that the accuracy fluctuation of TT-based option pricing is very small. We should note that TCI is a heuristic method, and depending on the choice of the initial points, the learning might not work well, that is, the error defined by Eq. (3) might not go below the threshold. Here, with the tolerance for TCI/SVD fixed, if the maximum relative error exceeds 4%, we redo TCI and SVD. The aforementioned mean and standard deviation are computed from successful cases.
@@ VI Summary and discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs. (10) and (12), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=9$, we demonstrated superiority in both computational complexity and time. For $d=10$ or $11$, the proposed method is comparable to the Monte Carlo method from the perspective of computational complexity. This is partly due to the need to adjust the integration mesh width $\eta$ for each asset, resulting in an increase in bond dimensions and, consequently, the computational complexity of contraction. However, we argue that if the computational complexity is comparable, TT-based option pricing holds an advantage in terms of computation time. This is because the TT-based method involves only multiplications and additions as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, iterated $dN_{\rm path}$ times. When the computational complexity of both methods was comparable, we confirmed that our proposed method was advantageous in terms of computation time. At this point, we also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths. However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\simeq 10$.
@@
Now, let us consider how the proposed method provides benefits in practical business in financial firms. An expected way to utilize this method is as follows. At night when the financial market is closed, we learn the TTs, and then, during the day when the market is open, we use these TTs to quickly price the option for the fluctuating parameters. If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial. In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process.
@@
Finally, we discuss future research directions. If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing. It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g. by increasing $N$ from the current value of 50 to, say, 100. It is a promising approach to incorporate parameter dependence using the Matrix Product Operators (MPO). It could eliminate the contraction between tensors related to parameters $p_{i}$ and $z_{i}$, potentially speeding up option pricing further. At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TT format is desired, although this study tried to incorporate either of them as the first step. Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.
@@ Acknowledgements.
We are grateful to Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to the code for tensor cross interpolation. R. S. is supported by the JSPS KAKENHI Grant No. 23KJ0295. R. S. thanks the Quantum Software Research Hub at Osaka University for the opportunity to participate in the study group on applications of tensor networks to option pricing. K. M. is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant no. JPMXS0120319794, JSPS KAKENHI Grant no. JP22K11924, and JST COI-NEXT Program Grant No. JPMJPF2014.

### v2

@@ V.2.1 The case of varying $\vec{\sigma}$
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi^{\phi}_{\mathrm{}}$ $\chi^{v}_{\mathrm{}}$
$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $3.53\times 10^{4}$ $5.0\times 10^{5}$ $2.47\times 10^{-4}$ $2.90\times 10^{-2}$ $4$ $6$
$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $6.38\times 10^{4}$ $6.0\times 10^{5}$ $5.61\times 10^{-4}$ $3.28\times 10^{-2}$ $5$ $6$
$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $1.00\times 10^{5}$ $7.0\times 10^{5}$ $6.80\times 10^{-4}$ $3.66\times 10^{-2}$ $6$ $7$
$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0258 0.0340 $1.54\times 10^{5}$ $8.0\times 10^{5}$ $9.71\times 10^{-4}$ $3.93\times 10^{-2}$ $7$ $7$
$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $1.99\times 10^{5}$ $9.0\times 10^{5}$ $1.37\times 10^{-3}$ $4.54\times 10^{-2}$ $7$ $7$
$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $3.00\times 10^{5}$ $1.0\times 10^{6}$ $1.88\times 10^{-3}$ $4.93\times 10^{-2}$ $7$ $8$
$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0377 0.0401 $7.33\times 10^{5}$ $1.1\times 10^{6}$ $6.25\times 10^{-3}$ $5.28\times 10^{-2}$ $9$ $14$
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{i}\in[0.15,0.25]$ and $S_{0}^{i}\in[90,120]$, and place 100 equally spaced grid points within these ranges. The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$. The term auto in (a) and (b) indicate that $\eta_{i}$ was automatically determined for each asset based on Eqs. (17) and (18), respectively. $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs. In TT-based option pricing with $\vec{S_{0}}$, the sharp increase in computational complexity at $d=11$ arises. This stems from the larger bond dimensions with a sufficiently low cutoff value for SVD in order to maintain its accuracy
@@ V.2.2 The case of varying $\vec{S_{0}}$
TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{i}$ and $z_{i+1}$ were reduced to around 10 (refer to Fig. 3(b)). However, at $d$ = 11, the computational complexity of TT-based calculation increases significantly. This is because it was necessary to increase the bond dimension by setting the SVD cutoff sufficiently low to maintain accuracy (refer to TABLE 2). This suggests that $d=11$ may already be reaching the upper limit of $d$ where accuracy can be maintained while surpassing the computational complexity over MC-based calculations.
Compared to varying $\sigma$, its accuracy is worse at $d=10$.
@@ V.2.3 Randomness in learning the TTs
Here, we mention the randomness in learning the TTs and the error induced by it. Since TCI starts with the randomly selected grid points, the accuracy of the TT-based method depends on these initial points. To assess such a fluctuation of the accuracy, for $d=10$, we evaluated the mean and standard deviation of the relative accuracy of the TT-based method in 20 runs, with initial points randomly selected in each. As a result, the mean was 0.0324 and the standard deviation was 0.0039, which indicates that the accuracy fluctuation of TT-based option pricing is very small. We should note that TCI is a heuristic method, and depending on the choice of the initial points, the learning might not work well. That is, the error defined by Eq. (3) might not go below the threshold. Here, with the tolerance for TCI/SVD fixed, if the maximum relative error exceeds 4%, we redo TCI and SVD. The aforementioned mean and standard deviation are computed from successful cases.
V.2.4 Total computational time for obtaining the MPOs
Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ through TCI and SVD. We focus on TCI because it dominates over SVD. It took about 3.4 hours for $d=11$ in the case that $\vec{\sigma}$ dependence is involved. This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec. VI. We also note that, when we do not incorporate the parameter dependence into the tensor networks for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence. In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1. This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
@@ VI Summary and discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs. (10) and (12), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=11$, we demonstrated superiority in both computational complexity and time. Here, we argue that, given at least comparable computational complexity, TT-based option pricing has an advantage in terms of computational time. This is because the TT-based method involves only multiplication and addition as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, to be iterated $dN_{\rm path}$ times. Indeed, in our demonstration, we confirmed that the TT-based method excels in computational complexity and is even more superior in terms of computational time. Note that both TT-based and MC-based option pricing implementations have room for improvement in computational time, and we consider the current comparison in time as reference data. We also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths. However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\geq 12$.
@@
Now, let us consider how the proposed method provides benefits in practical business in financial firms. An expected way to utilize this method is as follows. At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use these MPOs to quickly price the option for the fluctuating parameters. If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial. In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process and contraction.
@@
Finally, we discuss future research directions. If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing. It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g. by increasing $N$ from the current value of 50 to, say, 100. It is a promising approach to directly learning the MPOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing. At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TT format is desired, although this study tried to incorporate either of them as the first step. Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.
@@ Acknowledgements.
We are grateful to Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to the tensor cross interpolation code. We deeply appreciate Marc K. Ritter for his important remarks on the further improvement of our proposed method. R. S. is supported by the JSPS KAKENHI Grant No. 23KJ0295. R. S. thanks the Quantum Software Research Hub at Osaka University for the opportunity to participate in the study group on applications of tensor networks to option pricing. R.S. is grateful to Yusuke Himeoka, Yuta Mizuno, Wataru Mizukami, and Hiroshi Shinaoka, as the inspiration for this study was derived from their collaboration. K. M. is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant no. JPMXS0120319794, JSPS KAKENHI Grant no. JP22K11924, and JST COI-NEXT Program Grant No. JPMJPF2014.